what is the average speed if you hike 52 miles in 3 days
Found 2 solutions by solver91311, Alan3354:
Answer by solver91311(24713)



Average speed is distance divided by time. 3 days is 72 hours.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by Alan3354(69443)

d = r*t
r = d/t
r = 52/3 miles per day